ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 117, 2000

Aaronson, J. and H. Nakada Multiple recurrence of Markov shifts and other infinite measure preserving transformations

285

Abrams, L. The quantum Euler class and the quantum cohomology of the Grassmannians

335

Alon, N. Additive Latin transversals  125
Alspach, D. The dual of the Bourgain–Delbaen space  239

Bonet, J. and A. Defant The Levy–Steinitz rearrangement theorem for duals of metrizable spaces

131

Böröczky, Jr., K. The error of polytopal approximation with respect to the symmetric difference metric and the L_{p} metric

1

Camerlo, R. and A. S. Kechris Countable structures with a fixed group of automorphisms

105

D’Aquino, P. and A. Macintyre Non-standard finite fields over

IΔ_{0} + Ω_{1} 311

Defant, A. See J. Bonet

Eckmann, B. Introduction to ℓ_{2}-methods in topology: reduced ℓ_{2}-homology, harmonic chains, ℓ_{2}-Betti numbers

183

Farkas, D. R., C. Geiss, E. L. Green and E. N. Marcos Diagonalizable derivations of finite-dimensional algebras I

157

Feng, Q. Strongly Baire trees and a cofinal branch principle  71

Geiss, C. See D. R. Farkas

Giannopoulos, A. A. and V. D. Milman Extremal problems and isotropic positions of convex bodies

29

Green, E. L. See D. R. Farkas

Kanellopoulos, V. On the convexity of the weakly compact Chebyshev sets in Banach spaces

61

Kechris, A. S. See R. Camerlo

Kim, H. H. Langlands–Shahidi method and poles of automorphic L-functions II

261

Libman, A. Cardinality and nilpotency of localizations of groups and

G-modules 221

Macintyre, A. See P. D’Aquino

Marcos, E. N. See D. R. Farkas Milman, V. D. See A. A. Giannopoulos Nakada, H. See J. Aaronson Peres, Y., K. Simon and B. Solomyak Self-similar sets of zero

Hausdorff measure and positive packing measure

353

Simon, K. See Y. Peres Solomyak, B. See Y. Peres

Sun, H. On totally umbilical submanifolds of S^{n+p}  93